1- In the casting of steel under certain mold conditions, the mold constant in Chvorinov's Rule is 

known to be 4.0 min/cm2

, based on previous experience. The casting is a plate (fig. 1) whose 

length l = 30 cm, width w = 10 cm, and thickness/height h = 20 mm. Determine the time it will 

take for the casting to solidify.


1
Expert's answer
2021-05-26T16:11:01-0400

volume of the plate is

"V=30\times 10\times 2"

"V=600cm^2"

Area of the plate is

"A=2[(30\times10)+(30\times2)+10\times2)]"

"A=2(380)"

"A=760cm^2"


"T=c_m[\frac{V^2}{A}]"

"T=4\times[\frac{600^2}{780}]"

"T=4\times(0.789)^2"

"T=2.49 min"

therefore time required to solidify is "T=2.49 min"